Session: Nov 1997 (session 4)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Wed Nov 19 22:05:48 MET 1997) Success
appears to have ended succesfully

Onsala (A. Polatidis at Thu Nov 20 10:31:42 MET 1997) Success
No known problems. We recorded only LCP, expect signal on ODD VCs.

Effelsberg (dgraham at Fri Nov 21 09:46:44 MET 1997) Minor failures
Good weather. late on src at 0911.

Urumqi (Li Suqin at Fri Nov 21 11:44:49 MET 1997) Minor failures
It is cloudy. H-maser to GPS offset is 3.15us. Format to GPS offset is 3.03us. From beginning to 15:57:30 because the receiver trouble no data recored.

Jodrell (Alastair Gunn at Sat Nov 22 19:42:30 MET 1997) Success
Lost control of telescope briefly at about 11:00. No other known problems.

Metsahovi (Kaj Wiik at Mon Nov 24 13:38:26 MET 1997) Success
WX cloudy. All OK, some head positioning problems during tape changes, should not affect data.

Noto (C.Trigilio at Tue Nov 25 15:43:32 MET 1997) Success
Clear sky till 15 UT; raining at 16:30; OK at 21

Calibration (Fredrik T Rantakyro at Tue Dec 9 14:58:02 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name em027.antab Read the comments in the file to get further information.